# Limiting Rotation Rate of Neutron Stars from Crust Breaking and Gravitational Waves

### v1

@@ Abstract
Neutron stars are not observed to spin faster than about half their breakup rate. This limiting rotational frequency may be related to the strength of their crusts. As a star spins up from accretion, centrifugal forces stress the crust. We perform finite-element simulations of rotating neutron stars and find that the crust fails at rotation rates about half the breakup rate. We argue that this crust failure is asymmetric and produces a significant ellipticity (fractional difference in moments of inertia). A rotating star, with this ellipticity, radiates gravitational waves that limit further spin up. These stars may be promising sources for LIGO / VIRGO and next generation gravitational wave detectors.
@@ 1 Introduction
Neutron stars can gain angular momentum from accretion and spin rapidly Radhakrishnan & Srinivasan (1982). Indeed, there are many observed millisecond pulsars Manchester et al. (2005). However, no neutron star (NS) is observed to spin faster than about half the Keplerian breakup rate Hessels et al. (2006). Torques from gravitational wave (GW) radiation could limit the rotation rate Bildsten (1998). It is unclear what makes a NS asymmetric so that it radiates GW. Some possibilities include asymmetric electron capture reactions Bildsten (1998); Ushomirsky et al. (2000) or thermal expansion Jones & Hutchins (2024) in a star with an asymmetric temperature distribution. However, the produced asymmetry may not be large enough to explain the limiting spin Jones & Hutchins (2024). In addition to GW radiation, NS spin up may be limited by properties of the accretion disk-NS magnetosphere system Ushomirsky et al. (2000); Andersson et al. (2005); Haskell & Patruno (2011) or electromagnetic winds Parfrey et al. (2016). These effects may be too small to explain the limiting spin without GW radiation Çıkıntoğlu & Ekşi (2023).
There are many ongoing and near future searches for GW from rotating NS, see for example Riles (2023); Pagliaro et al. (2023). A number of these searches focused on the X-ray bright accreting system Scorpious X-1 Abbott et al. (2017a, b, 2019, 2022). These searches did not reach the sensitivity to detect a signal if one assumes torques from GW radiation and accretion balance each other. One search did reach a torque balance limit but only for some values of unknown parameters such as the spin frequency Zhang et al. (2021). No continuous GW signals have yet been detected. Future searches, with both existing and next generation detectors such as Cosmic Explorer Reitze et al. (2019), could be very promising.
@@
The rotation rate, when the crust breaks, depends on the strength of the crust. Many early works assumed the crust was weak and could break often. These breaks could produce star quakes and possibly explain some glitches in the rotation frequency of pulsars Baym & Pines (1971); Ruderman (1976, 1991). Molecular dynamics simulations find that the crust is very strong with a breaking strain (fractional deformation) of order 0.1 Horowitz & Kadau (2009). This strong crust may significantly change the picture. Crust breaking may now be rare. In an unpublished preprint, we estimated that the crust does not break until the rotation rate is high Fattoyev et al. (2018). In this letter we perform finite-element simulations to more accurately determine when centrifugal forces break the crust.
@@ 2 Formalism
We begin with a brief summary of our finite-element formalism Morales & Horowitz (2024) to obtain the structure of the crust of a Newtonian star with a polytropic equation of state. The first step consists of obtaining the zeroth-order stellar structure of a non-rotating, cold, barotropic, and Newtonian NS at mechanical equilibrium:
@@
Here $r$ is the radial coordinate, $\rho$ is the mass density, $p=K\rho^{2}$ is the pressure, $K$ is the polytropic proportionality constant, $\Phi$ is the gravitational potential, and $m$ is the enclosed mass. The prime () represents radial derivatives. When the star cools it crystallizes to form the solid crust with an angled-average shear modulus given by $\mu(\rho)\approx\kappa\rho$, where $\kappa=10^{16}$ cm2 s-2 Haskell et al. (2006). We assume that the bottom of the crust is located where the mass density is $\rho_{bottom}=2\times 10^{14}$ g/cm3 while the top of the crust is located where the mass density $\rho_{top}=10^{12}$ g/cm3. As we explained in our previous paper Morales & Horowitz (2024), placing the top of the crust at this high density does not affect the main results of the simulations and helps to avoid numerical difficulties.
@@
For each Legendre mode, the two boundary conditions that are needed to obtain $\delta U_{l}$ are the regularity of the perturbed gravitational potential at the origin and the interior-exterior matching of the perturbed gravitational potential. Since the latter boundary condition is implicit, the shooting method is used to solve for $\delta U_{l}$ in each l-th Legendre mode. With this scheme, the perturbed gravitational potential $\delta\Phi_{tot}$ is obtained. Then, we use that perturbed gravitational potential within the equations of hydro-elastic equilibrium to obtain the structure of the deformed crust using the finite-element method. Since the crustal mass is roughly 1 % of the star, we do not expect the crust to influence the perturbed gravitational potential significantly. This is the same approach that references Franco et al. (2000); Fattoyev et al. (2018) use. We use it in this work because it reduces the computational complexity of the finite-element method significantly.
@@
where : represents an inner product between two tensors. In the variational formulation, the displacement $\vec{u}$ is the trial function that we want to approximate while $\delta\vec{u}$ is the test function. $\overleftrightarrow{\varepsilon}$ is the elastic strain and $\overleftrightarrow{\sigma}$ is the sum of the elastic stress and the pressure. $\Omega_{s}$ is the integration domain (the solid crust) and $\partial\Omega_{s}$ is the sum of all the boundaries (the bottom and top of the crust). $\hat{n}$ represents a unit vector that is normal to the boundary surfaces (the bottom and top of the crust). Notice that the fourth term on the left-hand side of Eqn. (8) corresponds to the continuity of the radial traction across the bottom and top of the crust. The mesh to which we apply the finite-element method is an octant of a sphere with elements that have a size of approximately 100 m, see Fig. 1.
Figure 1: Mesh on one octant of the crust. The elements have an average size of 100 m, except near the bottom and top of the crust where the average mesh size is a few times smaller. In this plot, the mesh elements have been enlarged for clarity.
@@ 3 Results
The finite-element method yields the Eulerian displacements. These are shown in Fig. 2 where $u_{r}$ and $u_{\theta}$ represent the $r$ and $\theta$ components of the displacement in spherical coordinates. This is for a 1.4 $M_{\odot}$, $R=10$ km star that has been spun up from rest to $\Omega=220$ Hz. Material near the poles $\theta\approx 0$ moves in while material near the equator $\theta\approx\pi/2$ moves out to form the equatorial bulge.
Figure 2: Displacement $u_{r}$ versus $\theta$ (black). The thin black band shows $u_{r}$ is almost independent of $r$. $u_{\theta}$ versus $\theta$ near the bottom of the crust $R_{1}$ (gray) and near the top of the crust $R_{2}$ (light gray).
@@
The NS that spins up from rest to 220 Hz has the von-Mises strain field shown in Fig. 3. The von-Mises strain is largest at the base of the crust ($r=R_{1}$) and at the equator ($\theta=\pi/2$).
@@
We find that the crust breaks first at the equator of the star ($\theta=\pi/2$). This confirms the results that references Fattoyev et al. (2018) and Franco et al. (2000) obtained with their simple constant-density NS model. In addition, we find that the crust breaks first at the base ($r=R_{1}$). This assumes the breaking strain is independent of density.
Breaking at the base of the crust is in a high density region where nuclear pasta may be present. The breaking strain of pasta may be similar to that for the lower density crust Caplan et al. (2018), however elastic properties of pasta should be studied further.
Figure 4 shows that the maximum strain of the crust grows with rotational frequency. A strain of 0.1 is reached when the star is rotating at 0.6 of the Kepler breakup rate. This maximum strain is almost independent of the star’s mass. Results for a 2 M star are very similar to those for a 1.4 M star, assuming the same 10 km radius. Our simulations are for Newtonian gravity and a polytrope equation of state. This should be checked with relativistic calculations using more detailed equations of state. However, we do not expect large qualitative changes.
Perhaps the largest uncertainty is the exact value of the breaking strain. Molecular dynamics simulations predict that the breaking strain is of order 0.1. However the exact value could be somewhat different. For example a breaking strain of 0.05 is reached at 0.4 of the breakup speed. A major result of our simulations is that centrifugal forces break the crust when the rotation rate is about 1/2 of the breakup rate.
@@ 4 Discussion
When the crust breaks, the release of elastic stresses will allow parts of the crust to move and produce a nonzero ellipticity $\epsilon$. This is a measure of the asymmetry of the star,
@@
Here $I_{1},I_{2}$, and $I_{3}$ are the principle moments of inertia and the star is rotating about the 3 axis. The ellipticity is the important parameter to determine the characteristic strain amplitude $h_{0}$ of GW radiation from a star with rotational frequency $\Omega$ at a distance $d$ Riles (2023),
@@
We have not attempted to model in detail what happens to the crust as it fails and what the final crust configuration might be. Therefore, we can not directly calculate $\epsilon$. However, we believe $\epsilon$ is likely to be significant $\geq 10^{-8}$ because of the following considerations.
@@
    The maximum ellipticity the crust can support $\epsilon_{max}$ is large. If the whole crust is strained to near its breaking strain in such a way as to maximize $I_{1}-I_{2}$ (from a quadrupole shaped deformation) the ellipticity is large Morales & Horowitz (2022); Horowitz & Kadau (2009) ,
    $\epsilon_{max}\approx 10^{-5}\ .$ (14)
    This value is much larger that the $\epsilon\approx 10^{-8}$ that may be needed for torque balance.
@@
If $\epsilon\geq 10^{-8}$ after crust breaking, the system may rapidly reach torque equilibrium between GW radiation and accretion Çıkıntoğlu & Ekşi (2023) because the GW torque $\propto\Omega^{5}$ is a stiff function of the rotation rate. If $\epsilon$ is initially somewhat too large the star will spin down slightly from GW radiation until it reaches torque equilibrium. Alternatively if $\epsilon$ is somewhat too small the system may spin up slightly from accretion until it reaches equilibrium. In either case the system will reach torque equilibrium at a rotation rate near where the crust failed. Therefore, crust breaking may determine the maximum rotation rate of neutron stars.
We expect $\epsilon$ to evolve with time. Further crust failures, depending on where they are located, can either increase or decrease $\epsilon$. Elastic stress and $\epsilon$ can decrease with time due to viscoelastic creep Chugunov & Horowitz (2010). Further accretion can also decrease $\epsilon$ as crust material is buried to densities beyond the crust-core transition and melted. These effects could reduce $\epsilon$ to $\approx 10^{-9}$ after accretion stops. Woan et al argue that the observed spin down rates of millisecond pulsar populations suggest such a residual ellipticity Woan et al. (2018).
A population of rapidly rotating NS with significant ellipticity from crust breaking could be very promising sources for continuous GW searches Pagliaro et al. (2023); Reed et al. (2021). In general, rapidly rotating stars will produce stronger GW because of the factor of $\Omega^{2}$ in Eq. 13. Furthermore, search sensitivity will improve with improved techniques and observations and with better next generation detectors.
In conclusion, neutron stars (NS) are observed to spin no faster than about half the breakup rate. This may be related to the strength of their crusts. We performed finite-element simulations of rotating neutron stars and find the crust fails at rotation rates about half the breakup rate. This crust failure is likely to produce an asymmetric NS with a significant ellipticity (fractional difference in moments of inertia). Rapidly rotating stars with this ellipticity radiate gravitational waves that limit further spin up and may be promising sources for LIGO / VIRGO and next generation detectors.

### v2

@@ Abstract
Neutron stars are not observed to spin faster than about half their breakup rate. This limiting rotational frequency may be related to the strength of their crusts. As a star spins up from accretion, centrifugal forces stress the crust. We perform finite-element simulations of rotating neutron stars and find that the crust fails at rotation rates about half the breakup rate. Given uncertainties in microphysics, we have not determined the crust configuration after this failure. Instead, we argue that the crust may fail in an asymmetric way and could produce a configuration with a significant ellipticity (fractional difference in moments of inertia). If the ellipticity is large, a rotating star will radiate gravitational waves that may limit further spin up. These stars may be promising sources for LIGO / VIRGO and next generation gravitational wave detectors.
@@ 1 Introduction
Neutron stars can gain angular momentum from accretion and spin rapidly (Radhakrishnan & Srinivasan, 1982). Indeed, there are many observed millisecond pulsars (Manchester et al., 2005). However, no neutron star (NS) is observed to spin faster than about half the Keplerian breakup rate (Hessels et al., 2006). Torques from gravitational wave (GW) radiation could limit the rotation rate (Bildsten, 1998). It is unclear what makes a NS asymmetric so that it radiates GW. Some possibilities include asymmetric electron capture reactions (Bildsten, 1998; Ushomirsky et al., 2000) or thermal expansion (Jones & Hutchins, 2024). However, the produced asymmetry may not be large enough to explain the limiting spin (Jones & Hutchins, 2024). In addition to GW radiation, NS spin up may be limited by properties of the accretion disk-NS magnetosphere system (Ushomirsky et al., 2000; Andersson et al., 2005; Haskell & Patruno, 2011) or electromagnetic winds (Parfrey et al., 2016). These effects may be too small to explain the limiting spin without GW radiation (Çıkıntoğlu & Ekşi, 2023).
There are many ongoing and near future searches for GW from rotating NS, see for example (Riles, 2023; Pagliaro et al., 2023). A number of these searches focused on the X-ray bright accreting system Scorpius X-1 (Abbott et al., 2017a, b, 2019, 2022a). Many searches did not reach the sensitivity to detect a signal if one assumes torques from GW radiation and accretion balance each other. Recent searchs did reach a torque balance limit but only for some values of unknown parameters such as the spin frequency (Zhang et al., 2021; Abbott et al., 2022b). No continuous GW signals have yet been detected. Future searches, with both existing and next generation detectors such as Cosmic Explorer (Reitze et al., 2019), could be very promising.
@@
The rotation rate, when the crust breaks, depends on the strength of the crust. Many early works assumed the crust was weak and could break often. These breaks could produce star quakes and possibly explain some glitches in the rotation frequency of pulsars (Baym & Pines, 1971; Ruderman, 1976, 1991). Molecular dynamics simulations find that the crust is very strong with a breaking strain (fractional deformation) of order 0.1 (Horowitz & Kadau, 2009). This strong crust may significantly change the picture. Crust breaking may now be rare. In an unpublished preprint, we estimated, using a simple constant density model, that the crust does not break until the rotation rate is high (Fattoyev et al., 2018). In this letter we perform finite-element simulations, including a large range of crust densities, to more accurately determine when centrifugal forces break the crust.
@@ 2 Formalism
We begin with a brief summary of our finite-element formalism (Morales & Horowitz, 2024) to obtain the structure of the crust of a Newtonian star with a polytropic equation of state. [^1] Although we use the same formalism, Morales & Horowitz (2024) was focused on a different problem, how a small intrinsic asymmetry in the elastic properties of the crust produces an asymmetry in the shape of a NS. The first step consists of obtaining the zeroth-order stellar structure of a non-rotating, cold, barotropic, and Newtonian NS at mechanical equilibrium:
@@
Here $r$ is the radial coordinate, $\rho$ is the mass density, $p=K\rho^{2}$ is the pressure, $K$ is the polytropic proportionality constant, $\Phi$ is the gravitational potential, and $m$ is the enclosed mass. The prime () represents radial derivatives. When the star cools it crystallizes to form the solid crust with an angled-average shear modulus given by $\mu(\rho)\approx\kappa\rho$, where $\kappa=10^{16}$ cm2 s-2 (Haskell et al., 2006). For a fixed composition, the shear modulus of a Coulomb solid is $\propto\rho^{4/3}$. However given changes in composition with density, $\kappa\rho$ provides a reasonable fit. We do not expect much sensitivity to the exact form of $\mu(\rho)$. We assume that the bottom of the crust is located where the mass density is $\rho_{bottom}=2\times 10^{14}$ g/cm3 while the top of the crust is located where the mass density $\rho_{top}=10^{12}$ g/cm3. As we explained in our previous paper (Morales & Horowitz, 2024), placing the top of the crust at this high density does not affect the main results of the simulations and helps to avoid numerical difficulties.
@@
For each Legendre mode, the two boundary conditions that are needed to obtain $\delta U_{l}$ are the regularity of the perturbed gravitational potential at the origin and the interior-exterior matching of the perturbed gravitational potential. Since the latter boundary condition is implicit, the shooting method is used to solve for $\delta U_{l}$ in each l-th Legendre mode. With this scheme, the perturbed gravitational potential $\delta\Phi_{tot}$ is obtained. Then, we use that perturbed gravitational potential within the equations of hydro-elastic equilibrium to obtain the structure of the deformed crust using the finite-element method. Since the crustal mass is roughly 1 % of the star, we do not expect the crust to influence the perturbed gravitational potential significantly. This is the same approach that references (Franco et al., 2000; Fattoyev et al., 2018) use. We use it in this work because it reduces the computational complexity of the finite-element method significantly.
@@
where : represents an inner product between two tensors. In the variational formulation, the displacement $\vec{u}$ is the trial function that we want to approximate while $\delta\vec{u}$ is the test function. $\overleftrightarrow{\varepsilon}$ is the elastic strain and $\overleftrightarrow{\sigma}$ is the sum of the elastic stress and the pressure. $\Omega_{s}$ is the integration domain (the solid crust) and $\partial\Omega_{s}$ is the sum of all the boundaries (the bottom and top of the crust). $\hat{n}$ represents a unit vector that is normal to the boundary surfaces (the bottom and top of the crust). Notice that the fourth term on the left-hand side of Eqn. (8) corresponds to the continuity of the radial traction across the bottom and top of the crust. The mesh to which we apply the finite-element method is an octant of a sphere with elements that have a size of approximately 100 m (Morales & Horowitz, 2024).
@@ 3 Results
The finite-element method yields the Eulerian displacements. These are shown in Fig. 1 where $u_{r}$ and $u_{\theta}$ represent the $r$ and $\theta$ components of the displacement in spherical coordinates. This is for a 1.4 $M_{\odot}$, $R=10$ km star that has been spun up from rest to $\Omega=\Omega_{K}/2$ where $\Omega_{K}=(GM/R^{3})^{1/2}$ is the (nonrelativistic) Kepler frequency. Material near the poles $\theta\approx 0$ moves in while material near the equator $\theta\approx\pi/2$ moves out to form the equatorial bulge.
Figure 1: Displacement $u_{r}$ versus $\theta$ (black). This is almost independent of $r$. Also shown is $u_{\theta}$ versus $\theta$ near the bottom of the crust $R_{1}$ (gray) and near the top of the crust $R_{2}$ (light gray).
@@
The NS that spins up from rest to $\Omega_{K}/2$ has the von-Mises strain field shown in Fig. 2. The von-Mises strain is largest at the base of the crust ($r=R_{1}$) and at the equator ($\theta=\pi/2$).
@@
We find that the crust breaks first at the equator of the star. This confirms the results that Fattoyev et al. (2018) and Franco et al. (2000) obtained with their simple constant-density NS model. In addition, we find that the crust breaks first at the base ($r=R_{1}$). This assumes the breaking strain is independent of density. Breaking at the base of the crust is in a high density region where nuclear pasta may be present. The breaking strain of pasta may be similar to that for the lower density crust (Caplan et al., 2018), however elastic properties of pasta should be studied further.
Figure 3 shows that the maximum strain of the crust grows with rotational frequency. A strain of 0.1 is reached when the star is rotating at 0.58 of the Kepler breakup rate. This maximum strain is almost independent of the star’s mass. Results for a 2 M star are very similar to those for a 1.4 M star, assuming the same 10 km radius.
Our simulations are for Newtonian gravity. There are many fully relativistic calculations of the shape of rotating fluid stars, for example (Friedman et al., 1986; Konstantinou & Morsink, 2022). Perhaps the largest relativistic effect is a decrease of about one third from the non-relativistic Kepler frequency $\Omega_{K}=(GM/R^{3})^{1/2}$ to the relativistic Kepler frequency $\Omega^{r}_{K}$ (Konstantinou & Morsink, 2022), see Table 1. This reduction may also be due in part to an increase in the equatorial radius with increasing spin.
$R$ $\Omega_{K}$ $\Omega^{r}_{K}$ ${\Delta R_{e}/R}$ $\varepsilon^{\rm max}_{\rm vm}$
(km) (Hz) (Hz) 0.05 0.1 0.05 0.1
11 1890 1300 690 940 790 1100
12 1650 1130 600 810 690 960
13 1470 1000 530 720 620 850
Table 1: Rotational frequencies for a 1.4M neutron star of radius $R$. The non-relativistic (relativistic) Kepler frequency is $\Omega_{K}$ ($\Omega^{r}_{K})$ The columns $\Delta R_{e}/R=0.05(0.1)$ give the frequency (in Hz) where the equatorial radius of a fluid star increases by 5%(10%) (Konstantinou & Morsink, 2022). Finally, the last two columns give the frequency (in Hz) where the crust fails in our simulations for a breaking strain of 0.05 or 0.1.
We use a simple polytropic equation of state. Konstantinou & Morsink (2022) find nearly universal relations where the equation of state dependence is primarily through the star’s radius $R$ (via the $R^{-3/2}$ dependence of $\Omega_{K}$). Given NS radii determinations from X-ray and gravitational wave observations and nuclear experimental and theoretical information (Lim & Holt, 2022; Pang et al., 2023; Huth et al., 2022), we estimate equation of state uncertainties by considering $R$ from 11 to 13 km in Table 1. Perhaps the largest uncertainty is the exact value of the breaking strain. Molecular dynamics simulations predict that the breaking strain is of order 0.1. However the exact value could be somewhat different. For example a breaking strain of 0.05 (Baiko & Chugunov, 2018) is reached at 0.42 of the breakup speed.
Finally as a “sanity check” and to test for large relativistic effects we estimate crustal strain from relativistic calculations of the change in radius with spin of fluid stars. Qualitatively the crustal strain $\varepsilon_{rr}$ is of order $\Delta R_{e}/R$ where $\Delta R_{e}$ is the increase in equatorial radius with spin of a fluid star. Table 1 lists the spin frequencies where $R_{e}$ increases by 5% or 10% (Konstantinou & Morsink, 2022). These values are similar to but slightly smaller than our non-relativistic simulation results. This suggests that there is not a large error in our simulations and relativistic effects are not large.
Our simulations were performed for $R=10$ km. We scale our results to other radii by assuming $\Omega$ scales with $\Omega_{K}(R)$ because Konstantinou & Morsink (2022) find $\Delta R_{e}/R$ is approximately independent of $R$ for fixed $\Omega/\Omega_{K}^{r}$. This yields the $\Omega$ values in Table 1. For example, we find the crust fails at $\Omega/\Omega_{K}=0.42$ assuming the breaking strain is 0.05. The nonrelativistic Kepler frequency of a $R=12$ km star is 1650 Hz. Therefore, we predict that the crust of a $R=12$ km 1.4M star will fail at $\Omega=0.42\times 1690=690$ Hz. This frequency is comparable to the 716 Hz rotation frequency of the fastest known NS (Hessels et al., 2006). However, Table 1 indicates significant uncertainties.
@@ 4 Discussion
When the crust breaks, the release of elastic stresses will allow parts of the crust to move. This could produce a nonzero ellipticity $\epsilon$. This is a measure of the asymmetry of the star,
@@
Here $I_{1},I_{2}$, and $I_{3}$ are the principle moments of inertia and the star is rotating about the 3 axis. The ellipticity is the important parameter to determine the characteristic strain amplitude $h_{0}$ of GW radiation from a star with rotational frequency $\Omega$ at a distance $d$ (Riles, 2023),
@@
We have not attempted to model what happens to the crust as it fails and what the final crust configuration might be. This may require many assumptions about crust breaking microphysics. Therefore, we do not directly calculate $\epsilon$. However, we now argue $\epsilon$ is likely to be significant $\geq 10^{-8}$ because of the following considerations.
@@
    The maximum ellipticity the crust can support $\epsilon_{max}\approx 10^{-5}$(Morales & Horowitz, 2022; Horowitz & Kadau, 2009) is much larger that the $\epsilon\approx 10^{-8}$ that may be needed for torque balance. Before crust failure the deformations are axisymmetric and $\epsilon=0$. If crust failure releases strain on one side of the star, the remaining deformations may have a significant quadrupole pattern so that $\epsilon$ is nonzero and could be a noticeable fraction of $\epsilon_{max}$.
@@
If $\epsilon\geq 10^{-8}$ after crust breaking, the system may rapidly reach torque equilibrium between GW radiation and accretion (Çıkıntoğlu & Ekşi, 2023) because the GW torque $\propto\Omega^{5}$ is a stiff function of the rotation rate. If $\epsilon$ is initially somewhat too large the star will spin down slightly from GW radiation until it reaches torque equilibrium. Alternatively if $\epsilon$ is somewhat too small the system may spin up slightly from accretion until it reaches equilibrium. In either case the system will reach torque equilibrium at a rotation rate near where the crust failed. Therefore, crust breaking may determine the maximum rotation rate of neutron stars.
An accreting star can have a thick ocean and a magnetic field threading the crust that can respond to crust failure. Surface effects at low densities will likely not contribute significantly to $\epsilon$. However, they may produce electromagnetic signals associated with crust breaking, see for example (Tsang, 2013).
We expect $\epsilon$ to evolve with time. Further crust failures, depending on where they are located, can either increase or decrease $\epsilon$. Elastic stress and $\epsilon$ can decrease with time due to viscoelastic creep (Chugunov & Horowitz, 2010). Further accretion can also decrease $\epsilon$ as crust material is buried to densities beyond the crust-core transition and melted. These effects could reduce $\epsilon$ to $\approx 10^{-9}$ after accretion stops. Woan et al argue that the observed spin down rates of millisecond pulsar populations suggest such a residual ellipticity (Woan et al., 2018).
A population of rapidly rotating NS with significant ellipticity from crust breaking could be very promising sources for continuous GW searches (Pagliaro et al., 2023; Reed et al., 2021). In general, rapidly rotating stars will produce stronger GW because of the factor of $\Omega^{2}$ in Eq. 13. Furthermore, search sensitivity will improve with improved techniques and with better next generation detectors.
In conclusion, neutron stars (NS) are observed to spin no faster than about half the breakup rate. This may be related to the strength of their crusts. We performed finite-element simulations of rotating neutron stars and find the crust fails at rotation rates about half the breakup rate. This crust failure could produce an asymmetric NS with a significant ellipticity (fractional difference in moments of inertia). Rapidly rotating stars with this ellipticity radiate gravitational waves that could limit further spin up and may be promising sources for LIGO / VIRGO and next generation detectors.